OpenStudy (anonymous):

what is the rule to find the next term in the sequence? 7, 8.1, 9.2, 10.3, ...

OpenStudy (anonymous):

\[a _{n}=a _{n-1}+1.1\]

OpenStudy (amoodarya):

8.1-7=1.1 9.2-8.1=1.1 10.3-9.2=1.1 so an=an−1+1.1

OpenStudy (anonymous):

\[a _{1}=7\]